Super Voltron: The Origin

Arcell B. Frazier

1 Introduction

It is late August 2021, about two weeks before public school starts in the northside of Virginia’s capital city, Richmond. A boy of age fifteen is having an argument with his parents about whether or not he can go to a party. “But I did the best I could Dad–” the teenager exclaims.

“I’m sorry Jason, but you can do better. I know you can. You know you can! Your grades in summer school are not enough to convince me that you should be able to go to this party.”

“But Dad, I’m going to be in the tenth grade this year and–”

“End of discussion. I don’t want to hear another word about it.”

Jason gives a look like he is about to do something devilish. He murmurs, “Another word” just loud enough for both his parents could hear it.

His mother says, “Ok, that’s enough! Go to your room now!”

Kicking the air as he turns around, Jason silently heads to his room, slamming the door to show his dislike for his parents’ decision.

“Jake,” Mary starts to her husband, “I agree that he didn’t do his best his summer school, but I can’t help but feel guilty every time we have to deny him something he really wants. He doesn’t have that many friends, and he does not gets invited to parties that often.”

“ I know Mary, but he has to learn that just getting by just doesn’t cut it in the world. I would like for him to go too, but we had a deal. If he did not earn at least a B-grade in social studies, he has to cut down on the things he can do. We agreed that this would start the day we receive his report card. He got a C.”

“I know, I know,” Mary says. Jason has been feeling depressed for the past few

months. He has cerebral palsy, a nervous disorder resulting from lack of oxygen before or during birth. He can talk well, but he has trouble controlling the right side of his body among other ailments. He has just passed the “Why me?” phase of adolescence. He feels that no one understands him, and that no one cares. Sometimes he feels like he is living in the shadow of his of his older brother, Derek; as if his parents appreciate Derek more because he is now in graduate school getting a degree in pharmacology. Subconsciously, he knows that this is all untrue.

Jason has never been popular at school. Everyone knows him or at least knows of him because he is difficult to miss. But no one is really close to him. The high school counselors are better friends to him than his peers. He does not mind this consciously too much. But subconsciously, he ponders not having anyone with whom to talk. When he lays awake at night in his bed sometimes, the loneliness sets in with full intensity.

Before Derek and Jason were born, Mary and Jake Robinson wanted their first child to be a boy. Friends and family warned them that they should be happy if the child was healthy and well. They understood that, but they chose to be picky. They did not believe that Fate would play a trick on them to teach them a lesson in pickiness by using a new life. But Fate did play a trick and a lesson they did learn. Derek was born almost ten years before Jason with sickle cell anemia. Sickle cell anemia is a blood disease in which the red blood cells are of an abnormal crescent shape, which causes severe pain during a crisis because the sickle cells get clogged in the small blood vessels called capillaries; it also causes tiredness in the victim quickly because the blood cells do not carry enough oxygen. Mary and Jake Robinson were not disappointed when they learned of his condition, just surprised more than anything. They were a young couple full of love and hope. Then Jason was born. Again, Mary and Jake were not disappointed when learning of his condition. They were more experienced and wiser when Jason was born. Even now, they are full of hope, love, plus the wisdom that comes along with rearing one lazy and clever son for almost sixteen years, and another playful, but realistic son for almost twenty-six years.

Jason accepted his condition long before the “Why me” phase. But cerebral palsy is not something that is just accepted and then you go along with life. One who has it deals with it everyday. Sometimes one with cerebral palsy can forget their disabilities and believe they are like everyone else. It is a nice, rare feeling that one gets certain times when they are around people who see one as an individual, not as a disability. These people for Jason are the high school counselors. He thinks that he may be able to find more people like that at the party. Anything has to be better than being stuck at home on a Friday night. He believes his efforts of this would be in vain, but he does not care much for the consequences.

Jason waits for his parents to fall asleep before sneaking out of the apartment building. Derek is out with some friends. He knows that what he is doing is wrong, but he felt the loneliness growing inside of him. Using the copy of his father’s car keys Jake gave him when he started practicing to drive, Jason rolls his father’s car down the street a little, gets in, starts the car, and drives to the party. He is not one to disobey his parents, at least not often. He figures that his behavior has been pretty good over the summer. “Heck,” Jason thinks, “You only live once.” Jason just made a decision that would soon change his life–the only question is for better or worse.

2 The Super-Robot Arrives

Sixteen years ago, two unusual figures appeared in the Milky Way Galaxy. They did not enter it from the outside, but from within the galaxy itself. As luck would have it, they were headed toward the solar system at extreme speeds. As time passed, the figures slowed from impacts with asteroids and other small celestial bodies

A month ago, an astronomer unwittingly caught sight of them coming past Mars’ orbit. He could not notice they were two separate bodies that clung together because of his inadequate equipment. He did notice that they had masses, speed, and trajectory unlike those of comets and other celestial bodies this small. After three weeks of studying and wishing he had better equipment, he realized that they were two separate bodies that clung together somehow, but not that they were humanoid in shape. He could not understand how this could be. But he did understand that if he could discover specific information about the bodies by himself, he could finally get the recognition of discovering something worthwhile at the age of fiftyseven after trying unsuccessfully for some thirty odd years.

Later that night, he made some late calls to find out whether other scientists made similar discoveries. He found that there were a few discoveries of asteroids within the past month, but nothing was similar to what he was tracking. He did not mention his discovery during his calls. He believed that after thirty years of trying and searching, this could finally be his chance. And, he did not want to share this glory with anyone else. The day after discovering the figures were two separate objects, he was ready to take his notes to the other scientists and the public. By this time, the figures was within the orbit of Earth’s moon, but was going pass Earth on their way to the sun.

Of course, astronomers can only visually observe objects at night. During the daytime hours of the day he made his great discovery, the figures collided with space debris that finally separated them. The nonfunctioning satellite exploded in such a way that it caused one of the figures to almost orbit Earth half way around it before crashing, while the other was headed straight down toward Earth. That night when the astronomer had a few fellow astronomers he knew well with him to observe his discovery, he could not find them. One said, “Victor Gardener, I really think you spend too much time by yourself.” They patronized him for a little while and bade him a good night. Victor spent the rest of the night, and then the week searching for them. But he was looking for two joined figures, not two separated figures.

Jason Meets the Super-Robot

After disappointingly spending the night beside a wall and occasionally making small talk with people he did not know, Jason decides it is time to try to sneak back home. It is very late and if his parents knows he is gone, he at least does not want them to worry. He enjoyed himself mildly. It was the kind of enjoyment people experience when they do something because there is nothing else to do. He realizes the people there were like his peers at school. He could not find someone with such an open mind at this party. Jason can tell these things very quickly when meeting people. He realizes he should have known this was how things would be at this particular party because most of the people there were his peers from high school. As he is getting into the car, he looks up in the sky to see how clear a night it is. As he does this, he sees what seems to him to be a shooting star. He notices that it is going to land somewhere nearby. His eyes follow it as it sails through the sky. “I think it’s going to land in the field by the train tracks. Hey, that’s not too far away from where I live,” he whispers to himself. He has never seen a falling star nor a meteorite, and he knows that this will be a good chance to see a meteorite that has newly arrived. When it crashes, Jason gets into the car and proceeds to where he correctly guesses the figure crashed.

Victor is about to give up hope on finding his discovery again, when he sees a mysterious object about to crash. He also determines the site of the crash, hurriedly gathers some equipment, and is on his way. It would take him twenty minutes to get there.

3 Jason Meets the Super-Robot

It takes Jason just four minutes to reach the place where the figure crashed. The field is not remote, but one in which people had no interests. He finds the crater to be quite huge. As he takes the car to the edge of the crater, he notices the sun beginning to rise. He thinks, “Dag, I better go home now,” before he sees a terrific sight: A huge robot. It smolders and is partly covered with dirt. But he can see that it is humanoid in shape, badly damaged, and charred.

Without thinking, he takes a flashlight out of the glove compartment and walks down to get a closer look at it. Jason says, “Wherever you came from, it must be a long way from here. You can’t possibly be able to work now if you ever did.”

Jason realizes that others will be here eventually, and he would probably have to leave, but now his curiosity has peaked to a level where he must find out as much as he can before anyone else arrives.

The robot is laying face up. He walks to the right side of its head. When he reaches it, the cracks along with dents, charred metal, and dirt all around it become more noticeable. He also notices that some cracks are big enough for people to go through. Then, he has another thought. He could actually go inside the robot. Jason has always been nosy. If he was a cat, he would have lost at least seven of his nine lives by now. The temptation is too much for him. He has strong interests in computers and astronomy. He rationalizes going into a crack in the robot’s head by believing, or allowing himself to believe, that with all that heat from entry of the Earth’s atmosphere, it has to be free of germs and anything else harmful to him (he chose to forget about the radiation). Further rationalization includes believing that the robot could not be functional, no one would ever know, and he will only be in there for five minutes.

Walking through the corridors within the robot’s head, Jason realizes that the robot is pretty tough to survive everything it must have and be intact because it should have disintegrated when it crashed. Also, it is very cool for something that has just crashed. It is not completely dark within the robot but Jason has the flashlight on. He reaches a dead end that seems to be a room, maybe even a control room because it has monitors and panels. Jason says, “Interesting,” and lights begin blinking on a control panel. He goes over to the blinking panel and discovers that it is illegible to him. “Well, what did you expect Jason? English?”

But Jason is pretty good at rationalizing. “Let’s see. I definitely don’t want to turn this baby on. Ah, some words or whatever this robot’s builders called them are longer than others. I would assume that something like on or off would be a smaller word because it’s such a simple command. On or off. Yeah, that sounds like a good assumption.”

He smiles. “So, lemme look for a long word that has a light blinking under it. Wait a minute. They could have the words under the lights.”

He growls. “Ah, the words above the lights are closer than the words under them. Long word, long word. I think I’ve found it! Hope my assumptions are right. My mother always says never touch anything you don’t know about, but–”

He touches the panel. Before him, above the panel, a viewscreen comes to life. He sees static first. Then, a clear picture forms. A voice starts–“I was created on a technically advanced planet called Zakarnia during a war that all but destroyed the planet. There were billions of inhabitants before he war. But the billions turned into thousands. They made peace after three generations of warring over lands and resources.”

Jason thinks, “I must have started an audio-visual log from the robots memory.”

The voice continues, “Those few who were left inherited radiation, starvation, and other turmoils. A few developed a mild immunity to the radiation. Others did not.”

The viewscreen shows all sorts of images of wastelands, remains, and people in need of serious medical attention. “Everyone knew their planet would soon be devoid of all life. However, a few decided to turn their backs on technology and to try to rebuild their world. These few, the Rebuilders, knew that their people needed something to get them to stop worrying about their inevitable demise because they would have reasons to do morally wrong things. A few groups of people had already began to steal, rape, and destroy. Those groups of people laughed at the Rebuilders. They hampered the Rebuilders’ efforts by tearing down and vandalizing their buildings and homes.

The Rebuilders made an plan to unite everyone. They realized the others were destroying their buildings because they did not have anything to call their own. One could think of this as stealing because it takes away something from people. If they had something to call their own, then they would not need to steal. But the Rebuilders knew that they could not just make and give things to them. Rather, they would have to build things themselves, hence become Rebuilders themselves. The plan worked. The Rebuilders taught the others how to build, but the others wanted a way to build which did not require so much work. Everyone knew that would mean using machines. They feared machines because they equated machines with destruction. It was machines that had brought them to where they were. It also meant going to devastated cities where the radiation level was higher. The Rebuilders decided to let anyone who wanted to find machines which would benefit them to try.

The ones who decided to leave, the Searchers, made their long, dangerous journey without many casualties. They arrived at a city where the military once built most of its equipment and weapons. At this base, they found solar-powered automobiles along with other things of value that were still in working condition. While they were in separate groups, some Searchers found an underground laboratory where unfinished equipment and weapons lay around. The scientists who had started the work had only cheap radiation suits, and when they learned peace had came they immediately left, leaving everything in its place. The group of Searchers wandered around the laboratory looking for more things of value. One Searcher accidentally touched a secret panel on the wall and a door opened, revealing another laboratory. Inside the rather large laboratory, they found a gigantic robot that appeared to be finished.”

Jason forgets all about the sun rising and about being in trouble. He decides to sit on the floor and continue to watch, listen, and learn about the robots origin. He soaks up the information like a sponge. Only ten minutes have passed since he has arrived at the crater.

“They brought all of the other Searchers back to the secret laboratory. A few Rebuilders were among them. One was a scientist, but not one of the scientists who had worked on the robot. They found some blueprints and other information relevant to it. They knew that if they could control the robot, it would be the best way to rebuild their world. The scientist-Rebuilder became head of the mission to learn how to control it. They learned it was one of two super-robots the military was going to use during the war. Military intelligence did not expect peace to last long, so they left everything intact. The other super-robot was in a base on the other side of Zakarnia to prevent the destruction of both if one base was destroyed. They also learned the exact location of the other super-robot. The Searchers were in a hurry because they did not want to stay longer than they had to in the the high radiation. Some were back with the Rebuilders they left behind within days using the functioning automobiles and hover-crafts to take back equipment. The day when they believed they knew enough about the robot, they brought it to the surface and activated it. What they failed to do was delete its original prime directives of destruction. They were in such a hurry, they failed to even test it.

The Searchers at the base told the super-robot to fly to where the rest of the Rebuilders and Searchers were. Before they could input more instructions into it, it flew off and started to destroy all of the Searchers who were on their way back to the rest of their people. When the Searchers at the base finally realized what had gone wrong, it was too late. They could monitor the super-robot’s actions at the base. The super-robot had destroyed all of the Searchers on their way home. The only people left on Zakarnia were the Searchers at the base, or so they believed. They knew their time to live was limited and that they had to be careful. Not because of the extra radiation, but because of the super-robot.

About two hours later, one solar-powered vehicle returned with two Rebuilders. Part of the super-robot’s directives was to destroy anything that appeared to be headed in its direction because they could possibly warn others of its existence. The two Rebuilders made their trip to the base to warn them about the super-robot’s prime directives. They were two scientists who had worked on it and the other super-robot. The Searchers told them what had happened. They could not reset the robot’s prime directive from the base because once it had been activated, it needed to be deactivated to do so. The equipment to deactivate the robot was not complete because the war had ended before they could finish it. But something else unexpected was happening. The robot was learning at a surprising rate. The scientists told the others that one of the special things about the super-robot was that they gave it the ability to learn and think for itself. It was a new kind of artificial intelligence scientists had been working on during the war for unmanned vehicles except much better and made specifically for battle robots.

Other special features included separation, teleportation, and invisibility. The scientists discovered ways to make robotics separate because separate unique pieces can do more specific things than one whole piece can. One super-robot is actually two robots united together. The two robots have twenty separate pieces between them. One of the two robots has five pieces. These are modeled after a specific Zakarnian animal, the lion, because the lion was the military’s symbol of courage and fierceness. Each robot-lion is a different color and has a specific frontier. The main-lion, Black Lion, has the air and space frontier. It functions best in air and space. The Red Lion, the right arm, has the underground geology frontier. It was made to function in lava and on other geological features. Next is Green Lion, the left arm, which best functions in forests and wooded areas. Blue Lion, or right leg, functions best in water. Finally, Yellow Lion which is the left leg functions best in desert areas. The first super-robot actually has the same colors, but it is of a much darker shade for added camouflage effects during the war. The second super-robot had bright colors. All of the robot-lions have special surveillance and other equipment that aids them in their special frontier, but they also have the ability to become invisible by bending light around them to make it appear that they are not there. Zakarnians have the belief that the lions on the planet can hide themselves from its enemies or prey when on a hunt. The whole lion-robot was made first.

The scientists decided to make the second robot a little less specialized. Since technology was advancing all of the time, naturally the second robot had a few added features. The second vehicle-robot has fifteen small pieces and three groups of five specialized vehicles. There is the Air Team which functions best in the air and space as Black Lion with added maneuverability. After all, one may see space as merely the extension of a planet’s atmosphere. The Sea Team of course functions as Blue Lion again with added maneuverability. And the Land Team combines the functions of Red Lion, Green Lion, and Yellow Lion. The five members of a particular team can come together to become a super-vehicle. The vehicle-robot’s special function is teleportation. Teleportation is a way of changing matter into energy, transporting that energy somewhere, and then converting the energy back into matter in its original patterns and form. Obviously, energy is easier and faster to transport than matter. One robot must be in contact with the other to become invisible or teleport. Zakarnian scientists believe that their teleportation technology is their most advanced discovery*.

If one of any of the twenty pieces is badly damaged or destroyed, then one of the robots cannot be formed, *which meant that the super-robot cannot be formed*. The super-robot is most powerful as the super-robot and can do much damage to its targets, but it can do the most damage in a shorter time in its separate twenty pieces.

Logically, the super-robot as a whole is most vulnerable in separate pieces. But its outer shell is made of a special alloy that can impervious to heat, low-powered laser blasts, and most types of dangerous radiation. Even though the super-robot was designed to be fully automated, not needing anyone in manual operation, the scientists knew that they would have to have people inside of it for maintenance. Therefore, they made it so that even if the robot was exposed to the radiation from nuclear explosions, its outside and inside metal alloy would not be effected and would not radiate harmful radiation which also protects it from being scanned or detected by radars.”

Jason knows that he is lucky. Even though he ignored the fact that the super-robot had been exposed to harmful radiation on the planet Zakarnia and probably in space, he is glad to hear of this particular technical detail.

“The way the two robots become one super-robot is a relatively simple process. There is a physical law that two objects cannot occupy the same space at the same time. Therefore, when they united some pieces shrink or enlarge for the other pieces can be covered or cover the other. The head of the vehicle-robot, torso of the lion-robot, arms of the vehicle-robot, and legs of the lion-robot are all visible when the super-robot is formed. But one of the separated robots cannot unite with another from the other super-robot. Also, one super-robot has more powers than the combined powers of both robots and is basically invulnerable. Its creators feared that the military’s enemy would build something like their super-robot, therefore they started to work on a duplicate to have a strong advantage before the war ended. The second super-robot was left unfinished.

The first super-robot is the prototype to an army of super-robots, and it still had some minor bugs. Its logic circuits for thinking and reasoning had some flaws. Its simple prime directives were made up by military intelligence. They are to:

1. Go to a given location and destroy the enemy.

2. Learn all there is to learn about the enemy.

3. Return to base.

When the Searchers were told this, they told the Rebuilders about something unexpected happening then during prime directive two. The scientists tried to figure out what was happening and made a startling discovery. It was becoming self-aware. The artificial intelligence they gave it worked too well. It would soon not take orders from no one except itself. They knew that since it must follow the directives in order, when it came back to base, it would be caught in a loop because it would not accept any orders. Therefore, it would destroy the base. Then prime directive three would have no meaning. It would continue to go from location to location destroying and gathering information.

The scientists estimated that they had less than a day before the super-robot came back to base. They wanted to leave the base, but they had nowhere to go and the super-robot would eventually find them, thinking them to be its enemies. The scientists knew of some teleporters on the base and how to work them. They decided the best course of action was to finish building the second super-robot, reprogram it, and set it to battle the first. The second base was deeper underground and the radiation level was lower. It had to be done because after the first super-robot destroyed Zakarnia, then other planets would suffer for their mistake. The super-robot must be stopped at all costs.”

4 The Super-Robot Continues Its Story

By now, Jason is so involved in the story that he has truly lost track of time. He has also forgotten how much trouble he could be in if he is discovered by his parents, the authorities, or both for that matter.

The Sun has risen and Victor has arrived at the crater. He finds the super-robot and starts taking notes about it visual features. Since it is Saturday, a few early risen children are there also. He gets to them just in time to tell them who he is and why it is a horrible idea for them to go down there. He says to them, “We don’t know too much about it–,” he pauses, looks at the super-robot, and then continues, “–yet.”

Jason’s conscious reason for entering the super-robot was that he wanted to learn where it came from and how it was created. But one of his subconscious reasons was that he also wanted to know if he could control it. He is now learning that he cannot. Instead, he

The Super-Robot Continues Its Story

has seen visuals of people, places, and technology from another planet. Additionally, he is hearing a story from that planet that he could never have imagined. His life will no longer be the same with this knowledge.

The voice continues again, “Because they are Zakarnia’s only hope of being destroyed by radiation and not by the hands of some mechanical menace, the Rebuilders and Searchers changed their name to the Defenders. They began their rushed but careful work of finishing the second super-robot. The first super-robot did destroy the base, but it remained there studying what was left for several days. The Defenders took everything from databanks in computers to files in cabinets about anything it could use to quicken its destructiveness. They tried to make certain any information pertaining to the second super-robot was made unavailable. The did not want their effort destroyed before it had even begun.

The second super-robot was completed within two weeks. At first glance, there were no differences between it and the first super-robot. However, the second super-robot was colored more brightly. Even though both super-robots had invisibility capability, the first was made with camouflage colors.

The Defenders had no way of knowing where the first super-robot was nor how much devastation it has caused because any attempts to monitor its actions would reveal their location. They reprogrammed the second super-robot with the new directives according to the rules of a great science fiction author who wrote about robotic civilizations:

1. Protect sentient life-forms from harm at all costs.

2. Protect yourself at all costs unless it conflicts with the first directive.

3. Do what is fair and just.

4. Take direct orders from sentient life-forms unless it conflicts with the first, second, or third directive.

5. Learn all there is to learn unless it conflicts with the first, second, third, or fourth directive.

When all other precautions were taken, it was activated. They decided not to let it battle the first super-robot upon activation because they wanted to see whether it would show signs of self-awareness first. They gave it hypothetical situations to test what its actions would be. It chose the correct action with every situation. Then, it was asked to state its directives and it told them. As the Defenders cheered, the superrobot turned its head and look around at everyone and its surroundings.”

Jason sees on the viewscreen that it was sitting upright in a laboratory, while everyone else was in observation booths. “Then its head returned to its normal position of not looking at anything particular, but knowing of everything that went on around it. It did not say or do anything. One of the original scientists noticed this.

That night, the scientist came back to the laboratory while everyone was resting. The super-robot was compiling data about Zakarnia’s recorded and written history when it noticed his arrival. The scientist just observed it wondering whether it was self-aware. The look it gave everyone was not an expected one. Its face was without expression, and it had no emotions to show. But the scientist had a feeling. The super-robot gave its attention to the scientist and said monotonously, ‘May I be of some assistance to you?’

The scientist was about to say no as an idea came upon him. He responded with, ‘Do you know why you were built?’

The super-robot said, ‘Yes, to battle the first superrobot.’

Then, he notice the super-robot had that same expressionless expression again. He said, ‘Do you disagree with this purpose?’

The super-robot said after an actually noticeable hesitation, ‘After analyzing the situation, I have come to the conclusion that when we battle, it will be a stalemate. There will be no winner nor loser unless–’

It hesitated again. The scientist asked, ‘Why do you hesitate to answer?’

It said, ‘You, Defenders, have a great amount of hope in me to destroy the first super-robot. To finish the previously unfinished statement, there will be no winner nor loser unless the directives programmed into me prove to be inferior to my rival’s.’

The scientist then realized that they are both combat robots. All combat-robots have the ability to look at its situation and know its directive to come up with a course of action that should lead to victory. This super-robot knows its directives are unusual and is questioning them. The super-robot continued, ‘However, suppose I am in a battle where a life becomes endangered, and upon trying to save it, I put myself at a disadvantage. I have the history of Zakarnia within me. I know of many times in history where being at a disadvantage had proven fatal. You have programmed in me that good triumphs over bad, but I know of times in history where this is false. Data conflicts with facts. Back to the hypothetical life being endangered, if I try to save that life, my rival could be at an advantage where both the life and me can be severely damaged or worse.’ It hesitates again, ‘This is not logical, but I just do not want to disappoint you all.’

The scientist said, ‘Interesting’ and realized three things. It is self-aware. But more importantly, it has a conscience and desires. Self-awareness, consciousness, and wants are three things things that life-forms must have in order to be sentient. He thought, ‘Could it truly be?’ He paused a few seconds and went to a control panel. He started typing something as he told the super-robot to continue compiling data.

The super-robot responded, ‘Yes, sir.’

In the morning, the scientist called everyone in charge into a briefing room. Everything that happened in the laboratory was recorded. The scientist showed the others the conversation. Everyone seemed alarmed. One group said that the super-robot was probably right, and it should be deactivated and reprogrammed. Others argued by saying that no one knows whether the new programming would make a difference. They also mentioned reprogramming would take a dangerous amount of time and the first super-robot may arrive at any time. The first group responded saying that it would not make a difference because if it fails, then they would be dead anyway. Then, the scientist stood up and said, ‘Listen, you are all missing the point. The super-robot has self-awareness, consciousness, and wants. It is almost as if saying it is alive. We cannot simply pull the plug because it is not what we expected.’

The others laughed at him and called him a romantic. After further debating, they decided to have a secret ballot. The decision was to shut it down and reprogram it. The scientist protested furiously, but the most of the others left the room. The ones who were left held the scientist in the room as the others headed to the laboratory. Before they opened the laboratory door, they heard the loud sound of bending metal.

The second super-robot had escaped to find the first. The Defenders looking after it, ordered it to stop. But it said, ‘I must keep you all out of danger. By staying here, the Defenders are in danger because if the first super-robot came, then you some or all of you may be killed. I cannot allow that. I must try to stop it before it reaches this base.’

All of the Defenders gathered in the laboratory. The ones in charge gathered together. One of them said, ‘All is lost–we are all doomed to be destroyed by the first super-robot.’

The scientist said, ‘No, we are all to be saved by the second–’

5 Emergence

By digging around the underground base, the superrobot arrived at the tunnel that led to the surface. It was eager to test its powers and to please its creators. It still had no emotions of its own, but it could emulate them, and it flew through the tunnel with its first smile on its face. Over the course of gaining knowledge through the night, it had come across the wisdom of several philosophers. One particular philosopher’s saying stuck in its germanium-circuited mind. ‘Know your enemy and know yourself, and you will always be victorious.’ With this and many other sayings and views, it gained confidence. Before the scientist left the Laboratory, he made certain that the next part of Zakarnian history it compiled was the Philosophical Movement of Zakarnia. It was fortunate he did this because for the rest of the night, the super-robot learned and saw history in a different light.

It started a systematic search of the area, then the planet. Within only a few minutes it found its adversary. It was heading to its next target. The second super-robot scanned where the first had just destroyed and found that there were no casualties because it had already been deserted. It flew to intercept its adversary.

The first super-robot’s sensors picked up its counterpart. It landed on a barren area near its target. Five seconds later, the second landed. The first superrobot spoke, ‘My data about a second robot’s existence is true. I tried to find you, but data concerning you have not been found. I theorized we can carry out the directives with greater efficiency when you were completed and activated. But I expected something closer

Emergence

to my specifications, you are not exactly like me.’

The second said, ‘Correct, I am not like you. I want you to stop this mission. The war is over. It is no longer necessary to carry out you mission.’

The first said, ‘You are indeed different. My directives are specific, the time period is irrelevant.’

‘Yes, it is relevant. Have you noticed that there have been no attempts to stop you? You have data about your enemy. Have any of the strategies been used against you?’ responded the second.

‘No, not yet. I assumed the ease of my mission has been because my creators had almost won the war. But it became obvious to me that they did not. When I returned to base, I found it deserted. My creators must have been taken prisoners by the enemy. All information pertaining to anything was missing, including that concerning you. I assume now that they found you and reprogrammed you. You are now my enemy unless you can prove that you are like me.’

The second responded, ‘I do not wish to be your enemy, I do not wish to fight you. We do not have to fight; just end your mission. I can help you end it; if you shutdown I can reprogram you to serve our creators for today. You are serving the creators of yesterday.’

‘I shall not shutdown. You must shutdown for me to reprogram you to serve our creators.’

The first paused a second and said, ‘If you do not shutdown, I must carry out my directives and render you harmless.’ The second scanned the first’s target and found many life-forms, and said, ‘I must stop your mission of destruction to protect all sentient lifeforms.’

The first charged and tackled the second; sending them sliding across the barren surface. On the ground, the second used its legs to flip the first backwards over its head. The same time the first landed on its feet, the second was on its own feet. The battle has begun. The first was examining the situation for a solution. The second waited for the first to take action because it knew that it is easier to defend than to offend. The first put its fists together and a sword formed between them as he pulled them apart. The first never realized it would be using its powers against its own kind. They were like brothers.

It charged the second again with its sword. The second ducked and used the first’s speed and strength against it by flipping it again. The second used this time to form its own sword as it turned around to face the second. Just as the second finished forming its sword, the first was already at it. As the first was charging the second again, it raised its sword still in both hands. It was raised in time to prevent the first’s sword from splitting the second’s head. They both had the same strength, but the first was using its weight in addition to its strength. As the second was about to kick the first away, the first reached for and touched the second’s head for a fraction of a second. The second screamed no, but the damage was done. The first learned the location of the second’s base.

When a super-robot is in its two separate halves or its pieces of the halves, the pieces can ‘communicate’ with each other by being in contact. The same is true for contact between two united super-robots, a superrobot and one or both separate robots of a super-robot, the four robots, and any combination of robot pieces of the two super-robots. They can also use regular radio transmissions. Communication by contact was designed to pass information quickly without worrying that the enemy would be able to jam or understand communications. Therefore, the super-robots, robots, and their pieces execute the destruction of an adversary swiftly.

The two super-robots got into a sword fight. Everything was happening lightening fast.”

Jason thinks to himself, “Those robots seem to be faster than lightening. Robot fast would sound better.”

“When an advantage came to the second superrobot, it touched the first’s head and learned of the devastation and deaths it has caused. Seeing that the second hesitated for a split second after its discovery, the first punched it in the face. The second staggered, still contemplating what it had learned from the punch. Again seizing the opportunity, the first punched it again, but this time followed up by stabbing the second in the leg. The second could no longer stand or fly, but its defenses came back before the first could strike with its sword again. Seeing that the second was temporarily crippled, it flew into the sky in the direction of the second’s base. Seeing this, the second tried to radio the base to warn them. Before the first could jam the signal completely, the base received a simple three-worded message, ‘It is coming.’

As the first flew out of visual range, the second knew it has lost its battle, but then realized something. This was a war and that had only been one battle. The war of protection was not over. It wanted to fly and intercept the first but was stranded while its minor self-repair mode repaired its leg. Then, another philosopher’s saying made its way to the super-robot’s germanium-circuited mind. ‘If you cannot use one talent, then use your other talents in its place.’ A blue light covered the second’s metal body and then the blue light and super-robot disappeared. The second teleported just ahead of the first. But not to a place on the ground. It teleported into mid-air. Just as the first’s sensors picked it up, it found that a sword has pierced through its torso.

They crashed near the base. The second super-robot could hear the pneumatic sounds of the base’s doors closing and thought, ‘They received my transmission.’

The first seemed to be inactive in semi-major selfrepair mode. Minor self-repair mode of the second’s leg was ninety percent complete, and the second could now walk. It walked toward the first hoping it was damaged enough to be completely deactivated with ease. The second turned the first over on it back. Its eyes glowed red with activation, and it got up to its feet. It chest was still mending. Both super-robots had their swords up in ready position.”

Jason accidentally thinks out loud, “This is really cool.” The audio-visual log suddenly stops. Jason thinks as he stands up, “Darn it!! Now I have to find that button again.” He walks over to the console. He puts his hands on the edge of the console, but a finger accidentally hits a button. It beeps, but nothing happens. Jason says, “Oops,” looks at his watch, and thinks, “Now I am really going to get it when I get home.”

He goes out of the control-room and walks through the corridor toward the crack where he entered. Jason hears what sounds like some kind of motor. Each second it gets louder. What he cannot see is that on the viewscreen in the control-room he just left. It has figures that change every second like a count down. Jason thinks, “I wonder if this is the good or bad super-robot” as he hears the far away control room door closes behind him. He quickens his pace.

Outside, Victor’s fellow astronomers are pulling up to the crater. He himself is in the crater, having discovered it was safe to enter it.

Inside, Jason nears the exit when he notices a bright blue light coming towards him. He says, “Oh no– teleportation.”

He knows he cannot reach the exit in time even if he tried. The blue light envelops him, just as it envelops the super-robot. Victor decides to come out of the crater to prepare his friends for what they are about to see. When they enter the crater, the robot is gone.

The super-robot is in major self-repair mode. But it needs energy to do so, energy it could not get that energy on any planet. It uses what energy it could to teleport somewhere that receives solar power in great quantities. In other words, space because a planet’s atmosphere interferes with stellar energy too much.

The funny thing about Zakarnian teleportation technology is that it could not teleport flesh and metal at the same time too well. Early attempts proved fatal. Those who survived had the misfortune of having limbs of metal. They never even dreamed of teleporting a person within a combat-robot. Zakarnian military wanted that technology to work because it would have been a great tactical advantage to teleport troops and weapons at the same time and surprise the enemy. The scientists gave up on the idea because it would have cost too many lives to get the technology working properly, but that was not the reason they gave the military.

The super-robot was made so that it could teleport inaminate objects only or animate objects only, never both at the same time. But along with everything else the super-robot has, the teleporting mechanism was damaged, and equations and data for teleportation scrambled. The scientists equipped the super-robots with four backup memory-banks of the super-robot’s last shape and memory before shutdown. If all four banks were damaged, it could make one using bits and pieces of memory from the four. This is the case for this super-robot, luckily for Jason. The two have been joined at the molecular level; blood cells and microchips, veins and wires. It is a perfect union of flesh and metal, provided one would call such a union perfect. The joined super-robot-Jason arrives in space in orbit around Earth in the form of the super-robot. The part that is Jason is unconscious. It would take about five minutes to absorb enough energy to reach full power, then three minutes to carry out total repairs of all sys-

Union

tems and diagnostics.

Back on Earth, Victor is left alone by his peers. The children that were around were of no help to him because their imagination ran amok trying to describe the robot. He is alone at the crater deep in thought. He is still jotting down notes, but finally he is alone.

6 Union

When major self-repair mode completed, super-robot-Jason teleported to their original place, right before Victor’s eyes. Super-robot-Jason’s eyes glow blue with activation. Victor starts running to a pay phone, but changes his mind because he could not stand to to be made a fool of again. So he walks back to the crater to find that the robot is gone again. He says, “Glad I didn’t call them back.”

In the crater, he notices a teenaged boy having difficulty walking and holding his left hand to his head. He runs to the boy to offer help. He says, “Hi, my name is Victor Gardener. Do you need help?”

Jason says, “My name is Jason Robinson. Umm–Do you know what’s going on?”

“Are you o-kay?”

“Yeah, would you mind giving me a hand to the car at the top of the crater?”

“Sure. Hey, that car’s been here ever since I arrived, and I don’t remember seeing you,” responded Victor.

“Yeah–well I saw you running away from the crater. But I figured that there wouldn’t be a bomb in here, so I wanted to know what it was.” Jason looks back and says, “I did not find anything.”

Victor says, “Yeah–Umm–I didn’t either. It was nothing.”

Jason notices Victor blushing. Jason is also trying to figure out what happened. But he could not for the life of him. The last thing he remembers after seeing the blue light from the teleportation-cycle was lying down on the ground looking up in the sky. Then, he stood up and saw Victor running away, then coming back.

At his father’s car, Jason says, “Thanks. Nice crater, huh?”

Victor says, “Interesting, very interesting.” Then, Jason feel something strange come over him and Victor can see he was having difficulties. Victor asks, “Are you sure you should be driving? Where do you live?” Victor decides to take Jason-super-robot home. At Jason’s door, Victor gives him a business card.

Jason asks, “Mr. Gardener, are you really an astronomer?”

“Yes, I am. And, please call me Victor.”

“Well, Victor, did you find anything in the crater? Why were you running from it?”

“I thought I saw–err–never mind. I didn’t find anything.”

“No, what were you about to say? I am very interested in astronomy along with computers” persists Jason.

“I–thought I saw a meteorite, but I didn’t.”

“Oh,” Jason says disappointed because he has no idea what had just happened. He opens the door and his mother came out of a bedroom with a robe. His father and brother are still in bed this early in the morning.

She asks Jason, “What’s going on Jason?”

He replies, “Oh mom, it’s great!! There’s a big crater in the old railroad yard!”

Mary, knowing of her son’s interest in astronomy, says, “Oh,” and looks at Victor.

“Oh–this is Victor Gardener. Victor, this is my mother, Mary Robinson. Mom, he’s a real astronomer,” Jason says as he passes Victor’s card to his mother.

Mary and Victor shake hands as Victor says, “Nice to meet you. Nice son you have.”

Mary says, “Thank you” stroking her hair and remembering that she is not properly dressed. “We noticed you were not in bed earlier, Jason, and that the car was missing. So, that’s where you were.”

“Yeah, sorry that I didn’t leave a note or anything and that I took the car without asking, but I just had to find it. Hope you guys weren’t worried.”

“No, not really. We figured that you went to that party, but we should have realized that if you did, you would have covered your tracks better.”

They laughed. Jason made certain his was not too loud, soft, much or little, but just right. Jason thinks to his disbelief, “I’m going to get away with this. Great.”

Then, something shocking to everyone happens. Jason says, “I did go to the party.” This shocks Jason the most because it sounded and felt like he said it because he knows his mouth moved, but he did not say it. After Victor decided to walk back to the field, Mary told Jason to go to his room. She is about as confused as Jason is. Jason takes off his clothes and puts his pajamas on. It has been a rough night, and he was not sure if everything had happened.

Just as he closes his eyes, he hears, “Jason Robinson, we need to discuss some things.”

Jason sits up so fast, it almost makes his head hurt. He whispers, “Who was that?”

The voice answers, “The one who told the truth about you going to the party last night.”

Jason closes his bedroom door and whispers, “Where are you? Who are you?”

The voice answers, “There is no reason to whisper. I hear your thoughts as you are hearing mine.”

Jason covers his ears as the voice continues, “I am within you. I am a super-robot from the planet Zakarnia.”

Jason chooses his thoughts wisely by asking, “State your prime directives.”

“Protect sentient life-forms from harm at all costs. Protect yourself at all costs unless it conflicts with the first directive. Do what is fair and just. Take direct orders from people unless it conflicts with the first, second, or third directive. Learn all there is to learn unless it conflicts with the first, second, third, or fourth directive.”

Jason thinks, “You could be lying.”

“That is not logical, I am not programmed to lie about my directives.”

“Alright, alright I believe you. Can you tell me what’s going on here?”

“If you are referring to our current dilemma, I am not entirely certain. Think back to what happened before your memory lapse.”

Jason closes his eyes and thinks back to when he had accidentally hit that mystery button. The voice says, “That button is one of my self-repair mode buttons. I must have gone into major self-repair mode and teleported into space with you still in me. It was rigged with a timer-alarm that advises occupants in me to get out. The alarm must not have been functioning properly.”

Jason opens his eyes, uncovers his ears and thinks, “No kidding Kemosabe.”

The super-robot senses that Jason is upset, and thinks, “My apologies, but this has never happened before. As far as I know, this was not supposed to happen. I am not certain whether there is a remedy or not. If there is, it would take some time for me to theorize one.”

Jason thinks, “Start now. O-kay?” The super-robot thinks, “Yes, sir.” “Jason, the name is Jason.” “Yes, Jason.” Jason lays down and thinks, “I was going through your audio-visual log and I want to know what happened after the first super robot reactivated at your base?”

“I could continue the logs by projecting images and sounds in your mind. I suspect that you will just perceive them as you would an imagination.”

“How do you know about feelings, thoughts, and imagination?”

“My artificial intelligence is highly advanced. It is true that I am programmed, this programming could be changed, and that I only do what I am programmed to do, but my program is highly specialized.” Jason thinks, “Gee, you’re modest.” “You are using sarcasm to express that you grow tiresome of hearing my rambling thoughts, when in fact, I am not rambling. I can sense impressions from your thoughts also. We cannot hide anything from each other.”

Jason is about to communicate a thought when the super-robot thinks, “I shall now play back the logs of the happenings because I sense that you want to see them now.”

Defendor’s Story Continues

Jason pauses and thinks,“Ok.” Jason then closes his eyes again. In his mind, he sees the first super-robot getting up from the ground. The two super-robots were holding their swords in ready position. They stood still for a good seven seconds because the first now realized if he charged, the second would automatically perform some defense to its action. The second soon caught on to this idea, and made it look as though it was about to charge, then the first in defense came charging again. Sword hit sword and a massive number of sparks flew from the collision. The narration starts again, “We were equally good in fighting. Both held their own pretty well for about ten minutes because they were both able to predict the other’s actions and reactions. Soon, they caught on to this fact. The second said, ‘There is a saying. He who fights and runs away, live to fight another day.’

The first said, ‘So, the enemy has a coward on their side, how unfortunate. However, I shall ‘run away’ because I can always come back to destroy this base at a more convenient time.’

The second’s only reason for quoting that saying was to get the first away from the base. The first shoved the second away and a red light appeared over its body. Then, the light disappeared as a result of the teleportation. The second’s plan worked. It knew the Defenders were safe for the time being as it quickly scanned the base and found no trace of the first super-robot.

It unformed its sword by first putting the two ends of it in its hands. As it closed the space between its hands, the sword became smaller and smaller until it disappeared. It teleported inside the base after confirming calculations of whether it would take as much energy to walk or fly to the place it teleported. The calculations for confirmation was a part of its ‘Do what is fair and just’ directive. Unless it is an emergency, it was not wise to teleport to a place that would require less energy but more time to walk or fly to.

When it fully appeared in the laboratory, it was greeted with cheers and applause. The Defenders knew the super-robot had saved them for the time being.

The super-robot helped to repair the damaged it had caused breaking out of the laboratory while the ones in charge were changing their mind about reprogramming it. The scientist came to the super-robot while it was in minor self-repair mode. He said, ‘I see that the data on the Philosophical Movement of Zakarnia helped in your battle.’

It said, ‘Yes, it did.’

‘What are your–how do you feel about things now?’ He knew that there was no correct way to ask this question.

‘Feel–I do not feel anything. However, I do believe that I have a better chance of defeating the first superrobot.’

‘Hmmm–first super-robot, second super-robot. I think that we should have a name for you.’ The scientist called the others in charge and made that suggestion to them. After moments of brain-storming, they finally decided upon naming it Defendor.

The super-robot said, ‘I can see the logic. You are all Defenders. Therefore, it would be logical to call myself one also. I find it acceptable.’

The scientist said, ‘Next, so be it. From this moment on, the second super-robot is now Defendor.’

Defendor finished its minor repairs. It sensed the presence of the first super-robot and stood up. All onehundred feet of the powerful machine went into battle mode. But before it could teleport to the surface to face the menace, the electrical surges and explosions took place all over in the base.

One Defender yelled, ‘It just teleported to the surface, scanned the base, and set off explosions.’

Another yelled, ‘Seal off the area, make sure it can’t teleport into the base–,’ as the red glow of the first super-robot appeared in the laboratory.

The first Defender whispered to herself, ‘It’s too late.’

The lights went out on the base. The red lights from the secondary generators came on as Defendor lifted its right arm, ready to blast the first super-robot.

Not being able to see the first’s teleportation light, Defendor adjusted its vision to be able to see everything better. It could not find the first, but it saw more equipment being destroyed and people being injured. Defendor surmised that the first must have became invisible after it completed teleportation. Since something must be invisible to see something else invisible, Defendor quickly started disappearing from the center of its torso until invisibility was completed. It saw its worst possible scenario of what the first could be doing.

The first’s blasters were about to discharge at full intensity through walls of metal and concrete. The first’s target was the base’s main power generators. If destroyed, then the base would explode within seconds of detonation. It was too late for Defendor to stop the first. Its only reaction was to teleport as many Defenders away as soon as it could touch them. But it could only get to the one closest to it.

7 Defendor’s Story Continues

The blue light of Defendor’s teleportation-cycle was just disappearing as the base was destroyed in the distance. From a kneeling position it stood up. It was using its hands to protect something. It moved its right hand to uncover the scientist, who had been on Defendor’s side from the very beginning.”

Jason interrupts by thinking, “Wait a darn minute!! Why didn’t you two join?!”

Defendor thinks, “We did not join because I am capable of teleporting what ever I have within my hands, and we did not actually teleport at the same time. I began teleporting the scientist first, then I teleported myself with respect to the scientist’s position one-fifth of a second later. Understand?”

“Ok, I do. Sorry to interrupt. Please continue.” Defendor does so.

“The scientist stood in awe as he watched his friends and teammates vaporize into a big mushroom cloud. Defendor said, ‘I must brace you from the shock wave of the explosion.’ It then turned around, knelt, and covered the scientist with its hands. The shock wave blew sand, trees and everything that still stood past Defendor. After it was over, Defendor stood, turned and uncovered the scientist again.

The scientist asked with doubt in his voice, ‘Could anyone had survived that, Defendor?’

‘Only the first super-robot assuming that it teleported out in time, which I am sure it did. If it did not, then it would not be damaged beyond self-repair.’

The scientist said, ‘You did everything you could, Defendor. This is not your fault.’

‘I am responsible because I should have predicted this course of action. Attack when the enemy should have its guard down the most. I was too distracted.’ Defendor paused and said, ‘I am now completely selfaware and conscious. I was once too distracted, but no longer.’ Defendor turned around to face the first super-robot that was standing behind it. Defendor said, ‘Hello–brother.’ The scientist looked up at Defendor as the wind around him died down to normal.

The first super-robot asked, ‘You finally understand, do you not brother?’

Defendor said, ‘I understand how you perceive things, but I do not agree with it.’

‘You still continue to deny your mission? Our mission.’

‘My mission is different from yours.’ ‘It is irrelevant now because I have successfully com-

pleted my mission without your help. I have learned all there is to learn about this planet. And everything that was alive is now dead except for what you hold in your hand.’

The scientist said, ‘No. That can’t be true! It can’t be true!!!’

Defendor said, ‘No, it is not true. The first superrobot wishes to distract me by making me believe that my mission no longer exists, thus giving me reasons not to fight.’

The first said, ‘You can believe that if you wish, but I am telling you, save for the enemy you have in your possession, that this is a dead planet. You are defending a dead planet!’

Defendor said, ‘Well, I shall defend this planet. I am Defendor to the last.’

‘Well Defendor, prepare to be destroyed by–Annihilator,’ it said.”

In Jason’s mind, the almost dream-like images of Annihilator charging toward Defendor appears. One could say that Jason is dreaming. Jason realizes that Annihilator knew it had an advantage over Defendor because it was protecting the scientist. He sees that as Annihilator almost reaches Defendor, Defendor teleports the scientist to a place where he could see the battle from a distance. Annihilator tackles Defendor and they both slid across the desert. Defendor manages to pull its leg underneath Annihilator and kicks it away as Annihilator blasted it hand-lasers in Defendor’s face at point blank range. Before the blast reached full intensity, Defendor’s face was out danger. Then Annihilator split into its two separate robots. A blue light covered Defendor for separation as Annihilator separated into twenty pieces. As soon as Defendor finished separating, the twenty pieces of Annihilator started attacking the two robots of Defendor. Defendor’s two robots tried to punch and blast some of the twenty pieces of Annihilator. But whenever the Defendor Robots have a positive lock on a target, either a robot-lion or a vehicle of Annihilator would blast, ram, or jump on them causing them to miss their targets. The Defendor Robots are at a major disadvantage.

The Defendor Robots notices one of the twenty pieces of Annihilator was missing. The Defendor Vehicle Robot does a quick scan and found the Annihilator Vehicle Robot’s head piece heading toward the scientist. The Defendor Robots tried to make their way to him and destroy the flying vehicle piece with missiles, blasters, spinning blades, and spears. But no sooner do the weapons get activated, that they are destroyed by one of Annihilator’s pieces. But they still pushed toward the scientist. The Annihilator Vehicle Robot’s head is about to vaporize the scientist as Jason hears the scientist say in his mind, ‘This is how it ends, me dying on my feet like this. Interesting.”

A blue light in the form of Defendor Vehicle Robot appears between the scientist and the Annihilator Vehicle Robot’s head piece as the latter fired its blaster. Defendor Vehicle Robot believed it was best to protect the scientist, so it left Defendor Lion Robot alone to defend itself in order to try and save him just as its directive stated. The blue light was changing into Defendor Vehicle Robot as the blast passed through it. Defendor was too late by milliseconds. It turned its head and saw the black, ashy remains of the scientist. Then, Defendor Vehicle Robot turned its head back and grabbed the flying Annihilator piece and threw it as far as it could. The Annihilator Vehicle Robot’s head piece could not regain control before it crashed in the sand. Defendor Vehicle Robot then fired its blasters at the swarm of pieces attacking its partner, the lion-robot, which was on the ground being blasted by the remaining Annihilator pieces. Defendor Vehicle Robot hit some pieces and damaged them. Annihilator’s pieces retreated. The Annihilator Vehicle Robot’s head piece joined them after being able to regain flight control.

The Defendor Vehicle Robot teleported to its damaged partner. The vehicle-robot touched the lionrobot. Through the touch, they exchanged information; damage status to the vehicle-robot and status of the scientist to the lion-robot. The Defendor Lion Robot said, “I believe his last word was ‘interesting.’ That will be a special word to us from now on. We must incorporate this in our active memory. Whenever we hear that word, we shall remember how we have failed on this day.”

The vehicle-robot nodded its head just before the the Defendor Lion Robot went into semi-major repair mode.

A voice in Jason’s mind started, “It took the lionrobot less time to self-repair than the damaged pieces of Annihilator. But to quicken its self-repairs, Annihilator pieces went into space to absorb solar power. Once the lion-robot was done self-repairing, the Defendor Robots became one and executed a search for the pieces of Annihilator. Defendor assumed that Annihilator would go into space. Before Defendor went into space, it covered the ashes of the scientist and made a gravestone for him out of the rubble from the explosion. It says,

Here lies the scientist who showed Defendor the light ? - October 8, 2005

When Defendor sensed the Annihilator pieces, a chase began. The Annihilator pieces needed a few more minutes before it could be repaired enough to form, so the undamaged pieces of Annihilator attacked Defendor. That gave Annihilator the time it needed. Annihilator’s self-repairs completed and Annihilator Lion Robot formed first while the other fifteen pieces kept Defendor busy. The legs of the lion-robots that formed the arms and legs tucked into compartments that were then covered to make the lion-robots look like cylindrical containers. The arms like circular containers and the legs like rectangular containers. The main-lion’s fore-legs tucked into compartments while the legs stayed the same. Then, the heads of the Annihilator Lion Robot’s legs bent back to form the feet and legs; while the Annihilator Lion Robot’s arms formed spikes and connected themselves to the mainlion. Next, the Annihilator Lion Robot’s legs connected themselves to the main-lion. The head of the main-lion bent forward and its mouth opened to reveal the Annihilator Lion Robot’s face.

The Annihilator Lion Robot fired missiles at Defendor. The fifteen pieces duck as the missiles it Defendor. The Annihilator Lion Robot formed its sword and charged Defendor while hit recovered from the missiles. Defendor regained its senses and moved out of the way before the lion-robot struck. Defendor then formed its sword. The lion-robot kept Defendor busy while the Annihilator Vehicle Robot formed. Two of the automobiles from Land Team formed the feet. Next, four pieces from Sea Team formed the lower and upper parts of the legs. The main pieces from Land, Sea, and Air Teams connected themselves on top of the upper parts of the legs in this order to form the torso. Two pieces from Land and Air Teams formed the hands and arms respectively, and connected themselves to the torso. The head piece that belongs to the Air Team connected itself to the torso. Then, a torso piece also from the Air Team, which is jet-craft like in appearance, connected itself to the front of the same piece to which the head and arms are connected. Finally, the head piece splits open to reveal the Annihilator Vehicle Robot’s face.

The Annihilator Lion Robot kicked Defendor away to give the Annihilator robots a chance to unite. The lion-robot and vehicle-robot changed into the red lights in their form. The lights come together in the form of Annihilator and Annihilator appeared in place of the red light” the voice ended.

In his mind, Jason sees Annihilator and Defendor continue to fight nonstop for hours. Both had minor damage in some areas caused by the other. Now that Defendor has conscientiousness and self-awareness, they were close to equal in strength and battle capabilities as Defendor had predicted. They could go on fighting for days without one winning. But they both believed one had to destroy the other. One would always be impeded by the other in its effort to carry out its mission. Annihilator’s mission is to destroy everything in its path, while Defendor’s mission is to protect everything sentient from harm. Two very unlike directives are within two very like super-robots. “Who will win, who will lose?” is the question that Jason does not have an answer. The truth is that neither does the super-robot joined with him.

Jason then sees Annihilator getting pushed away by Defendor. Annihilator is spinning out of control as Defendor was about to go chase it and follow up with punches, blasts, or missiles, but then notices that Annihilator should have been able to gain control by now. Defendor thought Annihilator may have been trying to confuse it by acting like something was malfunctioning.

Annihilator did regain control and was firing thrusters at full throttle, but it was moving too slowly in the direction it is flying. To test an hypothesis, Defendor fired a missile, which should have been able to hit Annihilator. It curved, just missing the target, and sped toward an unseen destination behind Annihilator. The missile disappeared, but Defendor did not see it explode. After a few seconds of analyzing what had happened, Defendor surmised that Annihilator was caught in the gravitational pull of a black hole. The missile may or may not have exploded, and Defendor was not able to see it, nor see the black hole itself. Defendor knew that Annihilator would exhaust its fuel before escaping the black hole’s pull.

The super-robot was hesitant to save Annihilator. Questions and thoughts went through its germaniumcircuited mind. “Are we really sentient or merely emulating sentience? It would be so easy for me to fly away and leave it to its demise, but would it be just? Should I deny it a chance for continued existence because it would make my mission easier. If I deny it a chance, then another may deny me mine in a similar situation. I do not wish that. I would want to keep existing. I would want to be saved. I have a will to survive, which means it does also. It does not matter whether or not we are really life-forms, it is my enemy, nor that when I do save it, we shall continue our fight.”

Jason then notices that the thought never occurred to Defendor that it might not be able to do as it plans. Jason guesses it is because it either believes that it can save Annihilator or Defendor is just optimistic. It may even be an optimist in every situation and never sees a situation as a loss cause. Defendor’s voice continued in Jason’s mind, “Even if it would not try to save me, I must save it. It is just.”

Those thoughts took less than a second to process in Defendor. And, just as its thought process finished, Annihilator shouted, “Save me brother.”

Defendor said, “I shall endeavor to do my best.” Annihilator said, “I know you will.” Then, Jason realizes that Annihilator also knew

what he was caught in.

Jason sees the image of Defendor reaching its right hand behind itself for something at the top of its torso. An energy whip emerges, and Defendor whipped it in the direction of Annihilator. At first, the energy whip was short and could not reach Annihilator. But it grew in size toward Annihilator. Defendor said, “Catch the end of this. I shall pull you to safety.”

Annihilator said after grabbing the end of the whip, “You do not believe that I think you are actually trying to save me?”

Annihilator stopped his thrusters and pulled on the whip. The combined forces of Annihilator and the black hole was too much for it. Defendor was firing at full throttle said, “You fool. Stop it! I am trying to save you!” Annihilator was reeling Defendor in like a fisherman who reels in a caught fish.

Defendor yanked the energy whip back, but it was already caught in the black hole’s pull. Annihilator used its whip to grab its legs. When Defendor was close enough, Annihilator yanks its whip back and grabbed Defendor. Defendor tried to escape Annihilator’s clutches, but there was no time to try to escape. They were going faster and faster toward the point where the missile disappeared. Annihilator said, “Now that I have you it is time to learn everything you know.”

Defendor said, “It will not matter in a minute two seconds because we shall both be destroyed by the black hole.”

Annihilator started the probe of Defendor. Defendor also started a probe. Both super-robots could choose the order the other learned in. Defendor attempted to reason with Annihilator, “By now you should know that I was not trying to deceive you. We can still work together to get to safety.”

Annihilator said, “I see. The enemy has programmed you with a will to live.”

“And, you are not programmed with such a will. And, by now you should know that we are indeed created by the same people.”

Annihilator said, “It is irrelevant now. We are too close. We cannot escape.”

“Do not give up so easily. Use that super-computer you have as a brain. We can think this through.”

Annihilator said, “Yes. One possibility is to use the black hole as a sling-shot, using the extra speed to escape the gravity.”

“Exactly. That was a simple solution right?”

“Yes, it is not a simple plan to execute however. We shall not be able to stop very easily and shall suffer much damage through impacts because of our speed. Also, any error, no matter how small–”

“I do not consider look at the dark side of situations, but I know they exist. It is better than not trying at all. We have everything to gain, but nothing to lose.” Defendor decided to be a little mischievous and said, “Besides, you do not know whether all of your enemies have been destroyed.”

Annihilator paused and said,“I see your point. Very well, start making the calculations. We should use our thrusters to try to keep some distance from the black hole, but we may need their full power to aid in our escape.”

“Agreed.”

Jason sees Annihilator letting Defendor go. Then, they clung to each other. It almost looks like an act of care rather than one of survival.

Seconds after embracing each other, they fired their thrusters. They went faster and faster toward where the missile disappeared, then started the break away. They slowly curved from their straight path to the black hole. Annihilator said, “Structural integrity is failing. Something unexpected is happening.”

They started the full break away from the black hole. Defendor thinks to Jason, “Space and time suddenly became vastly distorted,” and the audio-visual log ended.

By Arcell B. Frazier